# Machine-Learned Potentials for Solvation Modeling

### v2

@@ IV.2.10   Delta-learned equivariant GNN for ionic media with implicit solvation
Coste et al. Coste et al. (2023) developed a deep implicit solvation (DIS) model for ionic aqueous systems using a $\Delta$-learned equivariant NN architecture. In this model, sodium and chloride ions are modeled explicitly, while water is treated implicitly, allowing accurate many-body potentials of mean force (PMFs) to be learned efficiently across varying salt concentrations (0.15–2.0 mol L-1). The ML model (Allegro) is trained via a force-matching loss against the difference between all-atom and a prior potential that models the van der Waals and electrostatic interactions, enabling accurate structural reproduction even in sparse ionic regimes.
@@
Summary: Table 7 summarizes case studies in MLP-aided solvation modeling, categorized by application type (A1 vs. A2), solvation paradigm (S1–S3), and MLP architecture (M1–M5). Two broad trends emerge. First, A1-type simulations, which involve learning the PES for configurational sampling, are predominantly coupled with explicit solvation (S1), especially in bulk or interfacial systems. In contrast, A2-type applications that target direct prediction of solvation-dependent properties—such as redox potentials or free energies—frequently adopt implicit solvation models (S2), reflecting their focus on energetic observables rather than full configurational sampling.
@@
IV.3   Sampling of configurational space
Efficient and accurate sampling of configurational space is a recurring theme across the development, training, and deployment of MLPs for solvation modeling. Case studies surveyed in Sections IV.1 and IV.2 illustrate how sampling-enhancing strategies—including active learning, metadynamics, replica exchange, and umbrella sampling—improve configurational diversity and ensure reliable prediction of solvation thermodynamics and reactivity.
Figure 12: Representative active learning or learning-on-the-fly strategies for MLP-based solvation modeling: (A) Generic active learning loop (adapted from Miksch et al. (2021)), (B) ML-accelerated MD workflow for catalytic processes (adapted from Chen, Zhang, and Zhang (2023)), (C) Energy-based structure selection for Diels–Alder reactions in water/methanol (adapted from Zhang, Juraskova, and Duarte (2024)), (D) C-NNP training at solid–liquid interfaces (adapted from Schran et al. (2021)). All subfigures are adapted from the original works for the purpose of scholarly synthesis and commentary. Original copyrights remain with the respective publishers.
Sampling-enhancing techniques—such as QbC Schran et al. (2021); Schran, Brezina, and Marsalek (2020), energy-based selection Zhang, Juraskova, and Duarte (2024), and extrapolation-grade filtering Chen, Zhang, and Zhang (2023)—are increasingly employed in both training and validation workflows. A recent contribution by Pan et al. Pan et al. (2025) demonstrates how pairing large-scale conformer enumeration with deep learning enables accurate tautomer ratio predictions in aqueous solution. Active learning or learning-on-the-fly (LOTF) is particularly useful for dynamically expanding datasets during MD simulations Csányi et al. (2004). Methods like DP-GEN Zhang et al. (2020), FLARE Yu et al. (2021), and AL4GAP Guo et al. (2023) rely on uncertainty estimation and targeted retraining to systematically refine force fields during simulations. C-NNP models Schran et al. (2021) and DeepPot workflows Célerse et al. (2024) exemplify successful LOTF integration.
Metadynamics, umbrella sampling, and replica exchange techniques bias sampling toward rare events such as hydrogen-bond rearrangements, proton transfer, and solvation shell exchange. The differentiability and low cost of MLPs make them ideal for coupling with these strategies. PIMD simulations further enable accurate modeling of NQEs Liu, Lan, and He (2022). MLPs can be integrated into umbrella sampling schemes, where configurations are sampled under harmonic restraints along collective variables (CVs) and later combined using weighted histogram analysis methods (WHAM) Kumar et al. (1992) or multistate Bennett acceptance ratio (MBAR) Bennett (1976); Shirts and Chodera (2008). When multiple system copies are simulated at varying conditions and swapped, replica exchange molecular dynamics (REMD) further improves ensemble convergence Fabregat et al. (2020); Sugita and Okamoto (1999). Accelerated MD techniques also benefit from MLP efficiency, and CV-based approaches expand configurational reach Goetz, Cruzeiro, and Roy (2020). A conceptually distinct approach is offered by Boltzmann generators Noé et al. (2019), which use invertible NNs to learn a direct mapping between a simple latent distribution (e.g., Gaussian) and the Boltzmann-weighted configuration space, showing promise for overcoming rare event bottlenecks in solvation modeling.
A recent study by Mendible-Barreto et al. Mendible-Barreto, Whitmer, and Colón (2025) offers a detailed analysis of how the choice and distribution of CVs in training datasets affect the ability of equivariant NN-MLPs to recover FES s. By systematically comparing training strategies based on classical and ab initio data for systems like butane and alanine dipeptide, the authors show that even sparse or biased CV sampling can yield accurate FES predictions if high-quality reference energies are used, thereby highlighting a practical route for sampling-aware MLP training.
Efficient and diverse sampling is essential not only during training data generation and uncertainty quantification, but also in achieving accurate predictions of solvation thermodynamics, reactivity, and interfacial behavior. These strategies enable broader configurational coverage and improved convergence of key observables in chemically and dynamically complex solvation environments. PIMD with DP-MP2 Liu, Lan, and He (2022) captures NQEs in RDFs and diffusion. However, non-conservative force-only MLPs may produce unphysical energy differences, biasing TI/FEP unless trained on both energy and force to ensure consistency: Inconsistent total energy, $U$, leads to unreliable free-energy, $F$, estimates as $F=U-TS$. Hence, conservative MLPs are essential for rigorous thermodynamic modeling.
Recent work by Axelrod and Gómez-Bombarelli Axelrod and Gómez-Bombarelli (2023) systematically examines the role of conformer ensembles in molecular property prediction using deep learning, emphasizing the importance of physically realistic conformer sampling, offering practical guidance for incorporating ensemble information into ML workflow. Global structure generators (e.g., ABCluster Elm (2020)) provide diverse initial states. Smith et al. Smith et al. (2018) used QbC to efficiently train ANI-1x using only 25% of the data, while Kulichenko et al. Kulichenko et al. (2024) emphasized uncertainty-guided selection. The importance of chemically meaningful and thermodynamically diverse configurations is further emphasized by the use of force diversity metrics Botu and Ramprasad (2015), electrostatic screening Behler and Csányi (2021), dimensionality reduction Friederich et al. (2021), and benchmark diagnostics Pinheiro et al. (2021). For spectroscopically sensitive solvation (e.g., hydrogen bonding), centroid-based self-consistent-charge density-functional tight-binding method (SCC-DFTB) has also been proposed Brezina, Beck, and Marsalek (2024).
@@ V   Open Challenges and Avenues
Electrostatic screening, dielectric response, and long-range charge transfer remain difficult to capture with standard local models. Periodic systems and interfaces introduce additional complexities such as spatial heterogeneity, solvent structuring, and directional noncovalent forces that challenge conventional descriptors and training protocols. For example, the molecular arrangement at the bulk is altered when liquids meet an interface Hayes, Warr, and Atkin (2015), introducing orientation-dependent polarization, and long-range structural correlations that are difficult to capture with local MLP descriptors. The TEA Challenge 2023 Poltavsky et al. (2025) highlighted significant model degradation under such conditions, especially for weakly interacting or rotationally mobile atoms (e.g., Pb in MAPbI3). Encoding physical priors such as many-body dispersion or nonlocal electrostatics is essential for accurate solvation-phase predictions. Equivariant architectures such as NequIP and MACE achieve high accuracy on training distributions, but often produce large maximum errors (up to 88 kcal mol-1 Å-1) in sparse or complex regions Poltavsky et al. (2025). Atom-wise force errors can vary 5–6$\times$ within a single molecule, complicating FF reliability. Unlike kernel methods (e.g., sGDML, GAP), current DNNs lack intrinsic uncertainty estimates, limiting their use in active learning and error-aware simulations.
@@
The practical deployment of MLPs is hampered by fragmented software ecosystems. Libraries for various NN-MLPs and kernel-models use different APIs and formats, complicating integration. Unified toolkits, pretrained model repositories, and user-friendly interfaces, such as ColabFit Kulichenko et al. (2024); Vita et al. (2023), can significantly lower adoption barriers. Consistent documentation, open benchmarks, and curated metadata for solvation systems are needed to support reproducibility and community development.
@@
Addressing these challenges will require coordinated advances in model architecture, training data design, hybrid physical–ML modeling, and community infrastructure. By improving the physical expressiveness, transferability, and usability of solvation-aware MLPs, future models can enable reliable simulations across chemical domains—from microsolvation to bulk liquids and complex interfacial systems.
@@ VI   Perspectives and Outlook
A notable gap in the current literature is the scarcity of workflows combining property-specific prediction tasks (A2) with hybrid solvation paradigms (S3); see Table 7. This absence may reflect both conceptual and technical challenges, for example, defining physically meaningful property labels across microsolvated clusters, or generating consistent training data from cluster–continuum models. These challenges are especially evident for solvent-sensitive properties such as valence and core-level excitations, NMR chemical shifts, or IR vibrational frequencies, where solvation induces spectral shifts that are difficult to decompose into local and bulk contributions Tripathi et al. (2025). Nevertheless, A2+S3 workflows represent a promising frontier. They can enable accurate, tractable predictions for systems where fully explicit solvation is prohibitive, yet pure continuum models oversimplify critical interactions. Future work using equivariant GNNs trained on curated microsolvation ensembles may bridge this gap and unlock new data-driven solvation strategies.
@@
Developing transferable, general-purpose MLPs for solvation is feasible only with chemically diverse datasets that span solute–solvent combinations and structural motifs. While small-molecule sets like QM9 offer a starting point, their solvation coverage is limited. Augmenting such datasets with varying numbers of explicit solvent molecules and ensuring solvent structure quality (e.g., via improved water models such as TIP4P/TIP5P) is essential. Moreover, capturing multipolar and anisotropic electrostatics via DCM or polarizable multipole potentials, such as the inexpensive atomic multipole optimized energetics for biomolecular applications (iAMOEBA) Wang et al. (2013) will be crucial for simulating dynamic environments. This addresses sampling sufficiency and solvent structuring questions, shedding light on bulk liquid structure—the neglected child, the “Cinderella” of condensed matter Elwell and Pointon (1972)—which, unlike gases and solids, remains far less well-characterized and understood.
The challenge of model transferability across solvents, temperatures, and compositions requires both architectural and data innovations. Transformer-based pretraining, generative models (e.g., diffusion autoencoders), and latent space alignment offer one route Zhang et al. (2025). Elm’s “clusteromics” concept Elm (2020) offers another route suggesting building structured databases of acid–base clusters and hybrid solvation structures, then unifying them through conservative MLPs and $\Delta$-ML corrections. This approach can scaffold general-purpose models for both microhydration and bulk solvation environments.

### v3

@@ IV.2.10  Delta-learned equivariant GNN for ionic media with implicit solvation
Coste et al. Coste et al. (2023) developed a deep implicit solvation (DIS) model for ionic aqueous systems using a $\Delta$-learned equivariant NN architecture. In this model, sodium and chloride ions are modeled explicitly, while water is treated implicitly, allowing accurate many-body potentials of mean force (PMFs) to be learned efficiently across varying salt concentrations (0.15–2.0 mol L-1). The ML model (Allegro) is trained using a force-matching loss to minimize the difference between all-atom and a prior potential, which models van der Waals and electrostatic interactions, enabling accurate structural reproduction even in sparse ionic regimes.
@@
Summary: Table 7 summarizes case studies in MLP-aided solvation modeling, categorized by application type (A1 vs. A2), solvation paradigm (S1–S3), and MLP architecture (M1–M5). Two broad trends emerge. First, A1-type simulations, which involve learning the PES for configurational sampling, are predominantly coupled with explicit solvation (S1), especially in bulk or interfacial systems. In contrast, A2-type applications that target direct prediction of solvation-dependent properties, such as redox potentials or free energies, frequently adopt implicit solvation models (S2), reflecting their focus on energetic observables rather than full configurational sampling.
@@ V  Open Challenges and Avenues
[…]
@@
The practical deployment of MLPs is hampered by fragmented software ecosystems. Libraries for various NN-MLPs and kernel models use different APIs and formats, complicating integration. Unified toolkits, pretrained model repositories, and user-friendly interfaces, such as ColabFit Kulichenko et al. (2024); Vita et al. (2023), can significantly lower adoption barriers. Consistent documentation, open benchmarks, and curated metadata for solvation systems are needed to support reproducibility and community development.
@@
The emerging “foundation” ML models that are pre-trained on large, diverse datasets and then fine-tuned on smaller, task-specific ones, may shape the next generation of transferable and data-efficient MLPs for solvation modeling. Similar efforts have already emerged as a prominent trend in the field of inorganic materials. Batatia et al. (2023); Lysogorskiy, Bochkarev, and Drautz (2025). Extending the foundation model paradigm to solvation modeling poses distinct challenges due to the dynamic, heterogeneous nature of liquid environments. Capturing solute-solvent interactions, long-range polarization, and composition-dependent effects will require large, chemically diverse liquid-phase datasets and careful fine-tuning strategies to maintain accuracy across solute-solvent combinations and thermodynamic conditions.
Addressing these challenges will require coordinated advances in model architecture, training data design, hybrid physical–ML modeling, and community infrastructure. By improving the physical expressiveness, transferability, and usability of solvation-aware MLPs, future models can enable reliable simulations across chemical domains ranging from microsolvation to bulk liquids and complex interfacial systems.
@@ VI  Perspectives and Outlook
A notable gap in the current literature is the scarcity of workflows combining property-specific prediction tasks (A2) with hybrid solvation paradigms (S3); see Table 7. This absence may reflect both conceptual and technical challenges, such as defining physically meaningful property labels across microsolvated clusters or generating consistent training data from cluster–continuum models. These challenges are especially evident for solvent-sensitive properties such as valence and core-level excitations, NMR chemical shifts, or IR vibrational frequencies, where solvation induces spectral shifts that are difficult to decompose into local and bulk contributions Tripathi et al. (2025). Nevertheless, A2+S3 workflows represent a promising frontier. They can enable accurate, tractable predictions for systems where fully explicit solvation is prohibitive, yet pure continuum models oversimplify critical interactions. Future work using equivariant GNNs trained on curated microsolvation ensembles may bridge this gap and unlock new data-driven solvation strategies.
@@
Developing transferable, general-purpose MLPs for solvation is feasible only with chemically diverse datasets that span solute–solvent combinations and structural motifs. While small-molecule sets like QM9 offer a starting point, their solvation coverage is limited. Augmenting such datasets with varying numbers of explicit solvent molecules and ensuring solvent structure quality (e.g., via improved water models such as TIP4P/TIP5P) is essential. Moreover, capturing multipolar and anisotropic electrostatics via distributed charge model (DCM) Devereux et al. (2014) or polarizable multipole potentials, such as the inexpensive atomic multipole optimized energetics for biomolecular applications (iAMOEBA) Wang et al. (2013), will be crucial for simulating dynamic environments. This addresses sampling sufficiency and solvent structuring questions, shedding light on bulk liquid structure—the neglected child, the “Cinderella” of condensed matter Tabor (1991)—which, unlike gases and solids, remains far less well-characterized and understood.
The challenge of model transferability across solvents, temperatures, and compositions requires both architectural and data innovations. Transformer-based pretraining, generative models (e.g., diffusion autoencoders), and latent space alignment offer one route Zhang et al. (2025). Elm’s “clusteromics” concept Elm (2020) offers another route suggesting building structured databases of acid–base clusters and hybrid solvation structures, then unifying them through conservative MLPs and $\Delta$-ML corrections. This approach can scaffold general-purpose models for both microhydration and bulk solvation environments. The advent of a foundation model for solvation could greatly reduce our reliance on manual, system-specific force field parameterization, paving the way for automated and reliable prediction of solvent effects even for newly designed or unconventional molecules.
@@
Acknowledgments
RB and SR thank the support of Mahindra University, under Project No. IRP/2024/24. SD and RR acknowledge the support of the Department of Atomic Energy, Government of India, under Project Identification No. RTI 4007.